OpenStudy (anonymous):

On Monday night, there were 69 unfilled seats in the theatre, and on Tuesday night there were 34 unfilled seats in the theatre. On Monday night, each ticket holder spent an average of $6.50 at the concession stand. On Tuesday night, each ticket holder spent an average of $4.00 at the concession stand. The profit made at the concession stand Monday night equaled the profit made Tuesday night. How many total seats are there in the theatre?

OpenStudy (callisto):

Let x be the total seats no. of seats filled on Mon.= x-69 no of seats filled on Tue. = x-34 Profit made on Mon = 6.5 (x-69) Profit made on Tue = 4(x-34) Since the profit are the same 6.5 (x-69) = 4 (x-34) Can you solve it from here?

OpenStudy (anonymous):

is it 125?

OpenStudy (callisto):

Yup, i think so

OpenStudy (anonymous):

great, thank you :D

OpenStudy (callisto):

Welcome :)